Excessive peaks with autotune.

Autotune performed on a 5977 GCMS. Autotune finishes to completion but is obviously not acceptable. Source cleaned and replaced with new filaments. Both filaments give similar results. Autotune file replaced with default. No differences observed.

GCMS instrument configuration performed.  Tune performed with similar bad results.  Instrument was used 3 weeks prior with no issues.  Electronic (board) issues?  Any suggestions appreciated.

  • This is a 5977A that was built in mid-2015. How many runs on it? .   High Energy Dynode (HED) related, Electron Multiplier (EM) related, signal related, or electronics related.

    The first thing to try is to go into manual tune and turn OFF the HED. I'd save the tune file with a new name like HED_OFF.U, so that the ATUNE.U file is not accidentally changed.

    The click on the "Scan" tab, turn on the PFTBA, then click on the "Scan" button, and look at the bouncing spectrum in the middle window on the left.

    With the HED off, it might require turning up the EM Voltage to get enough signal - a few hundred volts or so, maybe.  No peaks at all and it will say "no peaks" and turn the filament off again, so increase the EM volts a bit more and try again.

    Is the background noise gone but there's still signal/tune ion peaks?   It's either the HED itself, a plate with -10,000 volts on it that can be cleaned see: (+) Cleaning HED Detector Assemblies 5975-5977.pdf - Files - GC/MS - Agilent Community, the HED feedthrough O ring has a leak, the white HED cable is not connected properly or is faulty, or the HED power supply itself has failed.

    Is the background noise still there?  Does the noise go up and down when changing the EM Volts?   If so then suspect the EM horn has something in it, see the same HED cleaning PDF, the EM horn is bad, or there is an electronics issue.

    Or is it signal processing related? Stop scanning or profiling, turn off the PFTBA and filament, and remove the black coaxial cable from the side plate connector at the back of the analyzer door.  Don't be scared of the black wire it has very little to no voltage on it, unlike the white cable with -10,000V on it when it's turned on.

    In repeat profile mode, instead of Scan mode, click on the Profile tab and click the Profile button.  The top left window should show three plots of profile signal.  Is there noise? If not, then touch the end of the black cable with your fingertip.  Does signal respond somehow? If there's no noise and signal happens the data is getting from the EM to the electronics.

    Reconnect the black signal cable.

    Did those tests show something?





  • Thanks for the reply.  With the HED off and the system scanning and the voltage set to 2100, we were able to only see the 69 peak.  The white HED cable is properly connected.  Based on your previous suggestions, all things point to HED being the culprit.  I'll talk with management to see if they would like us to clean the plate or contact a service engineer.  Thanks again.  P.S we have had the system since 2015.  Cant really give you a good number of runs on it, but it can sit idle for many weeks in between runs.

  • A tiny bit of fluff on the HED or the hard wire that goes from the HED feedthrough to the HED can dance with -10,000 V on it.   The first easy thing to try is to vent, turn off, remove the two front cables that go to the analyzer, open the analyzer door very wide, and use an upright spray can of non-bitterant air duster or other ultra clean gas supply to blow out the entire area. Sometimes this is sufficient.    After that it would be following the posted cleaning document or getting a Field Service Engineer out to do it for you. Is the yearly PM due soon? - that extra cleaning could be done at that time.

Was this helpful?